OpenStudy (anonymous):

Need help with a math problem... 6x-36=y

OpenStudy (anonymous):

How would I solve this for x??

OpenStudy (anonymous):

\[6x-36=y\] add \(36\) to both sides and get \[6x=y+36\] then divide both sides by \(6\) to get \[x=\frac{y+36}{6}\]

OpenStudy (sweetburger):

solve for x would be 6x-36+y would become x=y/6+6

OpenStudy (sweetburger):

would you like a more in depth answer

OpenStudy (anonymous):

@TheRealMeeeee How petty that you sit behind a computer screen doing nothing but being a burden to others. Excuse yourself and make some real use of your time because you're wasting matter.

OpenStudy (anonymous):

which, as @sweetburger said, you can write as \[x=\frac{y}{6}+6\] if you likke

OpenStudy (sweetburger):

just a preference sorry

OpenStudy (sweetburger):

did our answers help?

OpenStudy (anonymous):

I got it now, thank you tons! :)

OpenStudy (isaiah.feynman):